# BibAgent: An Agentic Framework for Traceable Miscitation Detection in Scientific Literature

### v1

@@ 2 Taxonomy and MisciteBench Construction
We address this gap by introducing (i) a five-category taxonomy of miscitation errors governed by simple operational “litmus tests,” and (ii) MisciteBench, a contamination-controlled benchmark that instantiates this taxonomy at scale across all 254 Clarivate Journal Citation Reports (JCR) (Clarivate, 2025) subject categories and 21 broader disciplines. Conceptually, our taxonomy decomposes miscitation along five orthogonal dimensions—status of the source, factual content, scope of application, evidence strength, and attribution link—which jointly cover all failure modes observed in our corpus. MisciteBench realizes this taxonomy as a knowledge-blank, adversarial stress test for LLM-based citation reasoning, rather than merely a labeled dataset. Table 1 summarizes how legacy labels from prior miscitation studies are absorbed into these categories.
@@
2.1 A Unified 5-Category Taxonomy of Miscitation
Starting from a survey of existing typologies and a manual audit of hundreds of real-world miscitations sampled across all 21 high-level disciplines, we consolidate previously scattered subtypes into five conceptually distinct categories. Each category is defined by a diagnostic litmus question that makes annotation operational and reproducible.
@@ 2.1 A Unified 5-Category Taxonomy of Miscitation
These five categories were sufficient to label all 6,350 miscitation instances in MisciteBench (Section 2.2) without an “other” bucket. Annotators assign exactly one primary category per instance, guided by the litmus questions above. When multiple error types co-occur, we enforce a logical Dependency Precedence Rule that mirrors the verification process: checking a citation aborts at the first point of failure. The precedence order is:

### v2

@@ A.1 Unified 5-Category Taxonomy
Conceptually, our taxonomy decomposes miscitation along five orthogonal dimensions—status of the source, factual content, scope of application, evidence strength, and attribution link—which jointly cover all failure modes observed in our corpus. Table 3 summarizes how legacy labels from prior miscitation studies are absorbed into these categories.
@@
Starting from a survey of existing typologies and a manual audit of hundreds of real-world miscitations sampled across all 21 high-level disciplines, we consolidate previously scattered subtypes into five conceptually distinct categories. Each category is defined by a diagnostic litmus question that makes annotation operational and reproducible:
@@
These five categories were sufficient to label all 6,350 miscitation instances in MisciteBench without an “other” bucket. Annotators assign exactly one primary category per instance, guided by the litmus questions above. When multiple error types co-occur, we enforce a logical Dependency Precedence Rule that mirrors the verification process: checking a citation aborts at the first point of failure. The precedence order is